Calculate the volume in dm3 of oxygen evolved at s.t.p when a current of 5A is passed through acidified water for 198seconds. 1F = 96500C; Molar volume of gas = 22.4dm3]
1
Expert's answer
2019-03-27T04:36:01-0400

Anode:

"2H_2O(l) \rightarrow O_2(g) + 4H^+(aq) + 4e^-"

Cathode:

"2 \times (2H^+ + 2e^- \rightarrow H_2)"

Calculate the moles of electrons when a current of 5A is passed through acidified water for 198seconds:


"n = \frac{I\times t}{F} = \frac{5 A \times 198 s}{96500 \frac{C}{mol}} = 0.0103 mol e^-"

Calculate the moles of O2:

"0.0103 mol e^- (\frac{1 mol O_2}{4 mol e^-}) = 0.00256 mol"

Calculate vollume of O2:

"V = V_m \times n = 22.4 \frac{dm^3}{mol}\times 0.00256 mol = 0.0575 dm^3"

Answer: "0.0575 dm^3"